An AM radio station broadcasts an electromagnetic wave with a frequency of 554 kHz, whereas an FM station broadcasts an electromagnetic wave with a frequency of 96.4 MHz. How many AM photons are needed to have a total energy equal to that of one FM photon?
Answer #1

Here ,

let the number of the AM photons is N

energy of FM = N * energy of AM

h * (96.4 *10^6) = N * h * (554 *10^3)

(96.4 *10^6) = N * (554 *10^3)

solving for N

N = 174

174 AM photons are needed to have same energy as one FM photon